Reports of 17772
DGS (General)
satoru.ikeda - 22:24 Friday 28 January 2022 (19627) Print this report
Comment to Elapsed time output by FE dmesg is abnormal. (19527)

I restarted FE because the dmesg of k1ey0 and k1test0 were also having abnormal times.
After rebooting, normal time was output.

Non-image files attached to this comment
DGS (General)
satoru.ikeda - 17:29 Friday 28 January 2022 (19625) Print this report
Comment to Figuring out how to access the NFS server from Piconet (19528)

Ikeda, Oshino-san, Yamamoto-san

I changed the configuration so that I can access NFS (k1nfs0) from Piconet.
The LAN port is the third from the left of k1nfs0.

IP Address: 10.68.160.151
/export/kagra
/export/users
can be used when mounted.

The /kagra and /users were temporarily inaccessible from the DGS network, but we disconnected k1nfs0 from the network during this configuration.
They are now restored.

Non-image files attached to this comment
VIS (General)
satoru.ikeda - 17:19 Friday 28 January 2022 (19624) Print this report
Comment to Deletion of block with no mechanism in k1vispr{2,3,m}t (19473)

Change of VIS model file

Deleted the mechanicless part of BS.
The same contents as the previous K-Log#19256 were reflected to BS.

SRM is not a common file, so it is not included this time.

I did Revert(BS SDF),COEFF LOAD where there was a difference.

Git: 
Before: release/0.1.57
After: release/0.1.58
GpsTime:1327392621
 

Images attached to this comment
Non-image files attached to this comment
VIS (General)
satoru.ikeda - 17:02 Friday 28 January 2022 (19623) Print this report
Comment to Routine confirmation of DGS maintenance day (17831)

Routine check for DGS maintenance day (1/28)

Stepper motor
 [Checked]
  Check if the power supply can be controlled by BIO.
  Check if the script can be started.
  Confirmed that the limit switch installed is in Green.
 [Check result]
  No problem(BS was in use, so we'll skip it this time.)

Release Sensor
 [Checked]
 Confirmed that the Release Sensor is 0. 1: Seated
 [Check result]
 ETMX, ETMY, ITMX, ITMY : No problem

Picomotor
 [Checked]
  Ping Test is performed.
 [Check result]
  Known only

 NG: 
  MCI,STM1,IMMT2
  AS_WFS,POP_POM,POS_POM,REFL_WFS,POM1
  MCF(MCO,MCI)
   => Confirmation only, no recovery work was done.

 OK:
  ETMX,ETMY,ITMX,ITMY
  BS(IM,BF),SRM(IM,BF),SR2(IM,BF,STM),SR3(IM,BF)
  PRM(IM,BF),PR2(IM,BF),PR3(IM,BF)
  MCE,IMMT1,OMMT1,OMMT2,OSTM

  PCAL(EX1,EX2,EY1,EY2):Not confirmed because it is ON only when used.

Temperature Controller
 [Checked]
 Check if the temperature acquisition is working properly.

 [Check result]
  No problem

Precision Air Processer
 [Checked]
 Is the temperature acquisition function working properly?

 [Check result]
  No problem
 

Images attached to this comment
Non-image files attached to this comment
DGS (General)
takahiro.yamamoto - 15:59 Friday 28 January 2022 (19619) Print this report
DGS maintenance
Server maintenance
Security update was done for k1gate and gwdet.
These servers were rebooted a few times for this work.

Clean-up DQ channels
After O3GK, we stopped obsoleted guardian nodes.
But DQ channels related to a part of these nodes didn't removed.
So I removed them from K1EDCU_GRD.ini.

Old one was taken backup as K1EDCU_GRD.ini.klog19619.
I removed it in next maintenance day if there is no problem in next one week.
PEM (Center)
takahiro.yamamoto - 15:49 Friday 28 January 2022 (19622) Print this report
Comment to Update of PEM real time model (19618)

Use proper space for tests and don't save filters on an unrelated model!!


-----
I copied back chans/filter_archive/k1alsfib/K1ALSFIB_1325556129.txt which is latest and correct one as chans/K1ALSFIB.txt.
 

Images attached to this comment
VIS (BS)
fabian.arellano - 15:46 Friday 28 January 2022 (19620) Print this report
Calibration of F1 LVDT with remedying circuit.

With Hirata-san.

Summary: We calibrated the F1 LVDT, but only along 439 µm.

See pictures in the album BS Remedying work after O3.

Related klog entries:

  • Original calibration: 3350.
  • Remedying circuit: 18953.
  • Mounting of the displacement sensor: 18982.
  • F1 LVDT offset: 19580.

We used a new holder, which allows the adjustment of the sensor orientation in three degrees of freedom. This model of optical displacement sensor (Keyence LK-H022K) can only be used with specular reflection of light, so we had to tilt the sensor to the correct orientation (per the manual) and place a more reflective object on the BF cap (see ballast mass disc in the pictures).

Once  the sensor was giving a stable readout, we moved the F1 keystone with coil-magnet actuation. Although the sensor was working fine, it had a very limited displacement range of 439 µm. In order to use it along a larger range, we need to align it better.

The result of the calibration is m = 1.130 cnt/µm. The values to use in the medm screen are:

  • 1/ m = m_medm = 0.8848 µm / cnt.
  • offset_medm = -572 counts (this value produces a null output at 572 counts, per klog 19580).
Images attached to this report
Non-image files attached to this report
PEM (Center)
hirotaka.yuzurihara - 15:21 Friday 28 January 2022 (19621) Print this report
Comment to Update of PEM real time model (19618)

[Yokozawa, Yuzurihara]

We added the excitation signal on 'K1:PEM-MAG_EYC_BOOTH_EYC_X_EXC' from 15:00 ~ 15:15, by using awggui.

We injected the sine signal with 2, 5, 40, 110, 400Hz, and confirmed that the we can see the injected signal on the approriate BLRMS channel. (see the attached figures.)

Images attached to this comment
IOO (General)
kenta.tanaka - 14:48 Friday 28 January 2022 (19617) Print this report
Alignment the input axis for PMC to maximize the transmitted power

Miyakawa, Kenta

We aligned the input axis of PMC in order to maximize the transmitted power. As a result of alignment, the amount of transmitted light increased by about 10%. In addition, the camera could no longer see the higher-order modes. As far as the alignment monitor of the PSL was concerned, it was confirmed that the outgoing optical axis from the PSL had not changed.

### What we did

  • Before the alignment, TEM10 mode has been seen in PMC REFL camera (fig.1). From the shape of the beam seen by the camera, it was determined that the incident optical axis was shifted in the PIT direction. 
  • For safety reasons, we reduced the power of the beam incident on the PMC from 6 W to 2 W before the operation.
  • With the PMC locked, we adjusted the steering mirrors of the PMC (M2, M3 in KagraPslLayout.pdf) while observing the amount of light transmitted by the PMC (K1:LAS-POW_PMC_DC_INMON) and the shape of the beam using the REFL camera. 
  • As a result of alignment, the amount of transmitted light increased by about 10%. (fig.3) In addition, the camera could no longer see the higher-order modes. (fig.4)
  • As far as the alignment monitor of the PSL was concerned, it was confirmed that the outgoing optical axis from the PSL had not changed.
Images attached to this report
Non-image files attached to this report
PEM (Center)
hirotaka.yuzurihara - 14:31 Friday 28 January 2022 (19618) Print this report
Update of PEM real time model

[Yokozawa, Yuzurihara]

We added the BLRMS (band limited root mean square) channels for the magnetmeter, microphone, and accelerometer on PEM models (pemex0 and pemey0). You can see new channel names on attached text file.

The tail of channel name (such as `BLRMS_0MHZ30`) indicates the frequency band to compute the band limited RMS. In this case, the frequency from 0Hz to 30mHz is used.

Although the frequency for BLRMS of seismometer was up to 10Hz, these new channels are computed up to 500Hz.

Non-image files attached to this report
Comments to this report:
hirotaka.yuzurihara - 15:21 Friday 28 January 2022 (19621) Print this report

[Yokozawa, Yuzurihara]

We added the excitation signal on 'K1:PEM-MAG_EYC_BOOTH_EYC_X_EXC' from 15:00 ~ 15:15, by using awggui.

We injected the sine signal with 2, 5, 40, 110, 400Hz, and confirmed that the we can see the injected signal on the approriate BLRMS channel. (see the attached figures.)

Images attached to this comment
takahiro.yamamoto - 15:49 Friday 28 January 2022 (19622) Print this report

Use proper space for tests and don't save filters on an unrelated model!!


-----
I copied back chans/filter_archive/k1alsfib/K1ALSFIB_1325556129.txt which is latest and correct one as chans/K1ALSFIB.txt.
 

Images attached to this comment
FCL (Network)
shoichi.oshino - 14:08 Friday 28 January 2022 (19616) Print this report
Installed AC fan to the splice box at machine room
I installed AC fan to the splice box at the Y side of machine room.
After this installation work, the temperature inside the box becomes 25.3 degrees from over 40 degrees.
Images attached to this report
VIS (PR2)
takafumi.ushiba - 12:41 Friday 28 January 2022 (19614) Print this report
Comment to Suspension transfer function measurement and model construction: IM stage (19601)

I made models of PR2 IM stage at FM10 (SUSMod_220127) of IM_DAMP filter banks.
Figure 1 to 6 show the current models of IML, IMT, IMV, IMR, IMP, and IMY, respectively.

Then, I modified the gain and cut-off frequencies of 4th-order eliptic low-pass filters.
Since IMV and IMR has a strong resonance at the observation band, I set notch filters for these resonance at FM8 (notch) of IMV and IMR filter banks.

Theoretically new filters should work but I haven't checked yet because today is a maintainance day for DGS.
So, I wil check them after finishing today's maintanance.

Images attached to this comment
VIS (PR3)
takaaki.yokozawa - 12:10 Friday 28 January 2022 (19613) Print this report
Comment to Measurement of calibration factor for the length sensing oplev of PR3 (19578)
I calculated the calibration factor by inputting those length and angle;

const double d_mv = 833.0;
const double d_vb = 260.0;
const double d_bt = 50.0;
const double d_bl = 25.0;
const double d_lm = 230.0;
const double d_ml = 200.0;
const double alpha = 52.0;

const double meas_ver = 2.09;
const double meas_hori = -2.00;
const double meas_long = -7.82;

Then, we obtained
pitch=339.966019 [urad/cnt]
yaw=-218.722660 [urad/cnt]
long=-210.962107 [um/cnt]
LENS=310.555556 [1/mm]
VIS (General)
ryutaro.takahashi - 11:04 Friday 28 January 2022 (19612) Print this report
Comment to Limit of keystone position in LVDT count (18155)

The table of the limit in LVDT count for EY was updated. Blue cells indicate mechanical limit, red cells indicate FR limit, and yellow cell indicates out of LVDT range.

[EY]

  Min. Mid. Max.
F0 21200 22900 24800
F1 10700 3000 -4300
F2 -5500 3000 11800
F3 10400 2000 -7500
BF 5000 1880 -1240
DGS (General)
shoichi.oshino - 10:05 Friday 28 January 2022 (19611) Print this report
Kernel update on k1bck0
I updated the kernel on k1bck0 to the latest version.
I also update applications installed this server.
VAC (IX)
nobuhiro.kimura - 8:20 Friday 28 January 2022 (19610) Print this report
Comment to Result of Q-mass measurement at IX and EX (19590)

When: 27/Jan./2022
Who: Kimura and Yasui
Task:
We measured partial pressure of IXC and EXC by manual mode of Q-mass analzer.

Results are as follows;

  IX EX
Tortal Pressure (Pa) 3.1 x 10^-4 3.1 x 10^-4
He (Pa) 2.0 x 10^-10 7.7 x 10^-10
H2O (Pa) 7.5 x 10^-6 2.7 x 10^-4
O2 (Pa) 4.8 x 10^-8 1.7 x 10^-6
N2 (Pa) 2.1 x 10^-7 7.6 x 10^-6
Ratio of O2/N2 (%) 23 22

We confirmed main component of redisual gas in the cryostat was H2O.

We continue to measure and analyze redisual gas in the cryostat.

VAC (IX)
nobuhiro.kimura - 8:00 Friday 28 January 2022 (19609) Print this report
Comment to Result of the vacuum leak test at IXC (19426)
When: 27/Jan./2022
Who: Kimura, Yasui, M. Takahashi and Sawad (Hokuto)
Task:
We tried to find vacuum leak place in clean booth area of IXC.
Then, we confirmed there were no leaks.

We will continue vacuum leak test left area of IXC.
VIS (BS)
terrence.tsang - 4:31 Friday 28 January 2022 (19608) Print this report
BS Payload control and Guardian

Fig. 1-9 shows the IM and TM transfer function models.
Fig. 10-18 shows the OLTFs and the filters.

The controls for IMs are simple velocity damping filters with 4th order low-pass optimized for critical damping and 45 degrees phase margins.
The TM controls all come DC controls.

The TM yaw control was particularly problematic and there's a spurious ~46 mHz oscillation. This is also observable in the transfer function measurement in Fig. 8, which was being ignored during the fit.
I had to manually set the UGF of the integral control to 0.4 Hz so the gain at ~46 mHz becomes much higher. This however only suppresses the yaw peak-to-peak to around 2-4 urad (was something like 40 urad with my typical control filters).
The same oscillation occurs also in the IM stage.

The Guardian was tested and it worked bringing BS to the ALIGNED state.

Prior to the control work, the SENSALIGN matrix at the TM was also changed to minimize some cross-coupling.
K1:VIS-BS_TM_SENSALIGN:
[[ 0.99682099,  0.07718638, -0.04371401],
 [-0.04095595,  0.99441067, -0.01538625],
[-0.00125307,  0.03042447,  0.99889156]]

Images attached to this report
VIS (EY)
terrence.tsang - 3:20 Friday 28 January 2022 (19607) Print this report
Comment to ETMY troubles (19576)

With Dan Chen.

We've obtained a flash at the POS REFL PD.
Here's what I've done on the suspensions.

I've reworked F0-F2 GAS control and they worked fine now.
See Fig 1-6 for transfer function models and OLTFs.

The current MN to TM DC control doesn't work well with the 60 mHz oscillation for some reason and it actually amplifies the oscillation.
As a temporary solution, I reduced the gain by 10 times, moving back the UGF from 0.1 Hz to 0.01 Hz.
I've also put a notch at 60 mHz, hoping to reduce the interaction between the 60 mHz and the DC control.
When the optical lever is in range, the MN to DC control works well for steering the mirror in yaw but the oscillation is persistent.
It naturally damps out so it's likely not excited by some external source. But it takes hours for it to reduce to something say 0.1 peak-to-peak normalized QPD readout (assuming ~100 urad/count, this converts to 10 urad peak-to-peak).
I have a feeling that this mode was excited by large yaw adjustment, like the transient of the BF yaw DC control, or BF yaw pico motor control.
This situation is unavoidable since the BF yaw is currently far from "zero", sitting at -2600 urad when it's not controlled. And, it couldn't be offloaded using the F0 yaw stepper motor (we might try again later).
Currently, I don't think there's an easy solution for this 60 mHz mode for now if it ever gets excited again except using an external optical lever at the MN stage, like what was done in IX.
Anyway, this oscillation may be acceptable for PCal work so we may be able to proceed.

With all controls manually engaged, we were able to steer the ETMY.
We first checked if the old optical lever reference coarsely aligned the reflection beam.
We again made an iris for the duct at EYA using aluminium foil and moved the ETMY in yaw. We did see a reflected beam moving horizontally so the optical lever reference was not bad.
We move the ETMY back to the original position and removed the foil, but we can't see any changes in the POS REFL PD.
We tried scanning ETMY in pitch and yaw but still nothing.
Then we decided to go back to the center area and look for the reflected beam but we couldn't identify the reflected beam from ETMY.
After that, we tried to scan ETMY in yaw by large amount again and with pure luck, we saw a single flash.
We then identify a good yaw position and put it to the set point. We tried to optimize the pitch as well but it's difficult as the 60 mHz keeps misaligning yaw and we can only see flashes.
In the end, we settled on a setpoint (SDF shown in Fig. 7)

K1:VIS-ETMY_TM_SET_P_OFFSET: 0.15
K1:VIS-ETMY_TM_SET_Y_OFFSET: -0.56

Fig. 8 Shows the flashes with the TM oplev readout. (note the REFL output sits at ~0.6 and goes up to ~0.8 when the ETMY is aligned)

Images attached to this comment
VIS (SR3)
terrence.tsang - 2:39 Friday 28 January 2022 (19606) Print this report
SR3 TM setpoint and IM OPTICALIGN

During the green Y alignment work, we changed the setpoint of SR3 to

K1:VIS-SR3_TM_SET_P_OFFSET: 12
K1:VIS-SR3_TM_SET_Y_OFFSET: -2

Also, a few days ago I discovered that the TM DC control doesn't have enough actuation range to put the TM to setpoint so I have added some -250 counts offset at K1:VIS-SR3_IM_OPTICALIGN_Y_OFFSET.

VIS (PR3)
fabian.arellano - 21:04 Thursday 27 January 2022 (19604) Print this report
Comment to In-vacuum cable health check at outer side of flanges. (18108)

With Hirata-san.

Summary: We checked the cable problems reported in klog 19501. All the cables of the suspension seem to be in good condition now.

  • BF V3 LVDT: the problem was solved on the 7th of October 2021 and reported in klog 18497, so we didn't examine the cable.
  • Traverser R2 cable: we measured again the resistances that were too high, and we obtained the expected values. Likely, there was an error in measuring them before, so we didn't do anything else.
  • Traverser R1 cable: We measured again the resistances that were too high. For some of them, we got the expected values, but for some others, we corroborated the values were too high. We identified the origin of the malfunction in the in-vacuum connector that connects to the flange. Touching the cables changed some resistance values by a lot. We removed the cable and Hirata-san crimped all the pins again. This solved the problem.

I attached the compilation of the related information of all PR suspensions with the updated values for PR3.

Non-image files attached to this comment
VIS (General)
ryutaro.takahashi - 20:59 Thursday 27 January 2022 (19603) Print this report
Noise level of new accelerometers

The noise level of installed accelerometers is higher than the level expected from the prototype test. So we are investigating the causes.

I measured the spectrum of the accelerometers on the IP in IY, and compared them (FLDACCINF) with the cases of the LVDT driver off (cable connected) and the cables disconnected to the ADC. Though the noise level at low frequencies (<0.1Hz) is smaller than these cases by servo gain (2~5) due to the servo loop, the noise level was comparable with the ADC noise.

Images attached to this report
FCL (Air)
shinji.miyoki - 20:06 Thursday 27 January 2022 (19600) Print this report
Tunnel between BS and PR area

[Hayakawa, Takahashi.M, Oomae]

A clean tunnel between PR and BS area. An additional FFU might be added on this tunnel. 

I am now asking  Hayakawa-kun how to have stuff go through this tunnel.

Images attached to this report
VIS (PRM)
takafumi.ushiba - 20:00 Thursday 27 January 2022 (19602) Print this report
Suspension transfer function measurement

I measured the TFs of BF stage and GAS flters of PRM.
Measured data was stored at /users/VISsvn/TypeBp/PR2/TF/Measurements/20220127/.

I will make models based on the measurement at some point.

VIS (PR2)
takafumi.ushiba - 19:57 Thursday 27 January 2022 (19601) Print this report
Suspension transfer function measurement and model construction: IM stage

Similar work with klog19597 and klog19598.

I measured the TFs of IM stage.
Measured data was stored at /users/VISsvn/TypeBp/PR2/TF/Measurements/20220127/PLANT_PR2_STANDBY_IM_TEST_{L,T,V,R,P,Y}_20220127.xml.

I haven't finished making models of L, T, and V, but not yet finished to make those of R, P, and Y.
So, I will continue to make model.

Comments to this report:
takafumi.ushiba - 12:41 Friday 28 January 2022 (19614) Print this report

I made models of PR2 IM stage at FM10 (SUSMod_220127) of IM_DAMP filter banks.
Figure 1 to 6 show the current models of IML, IMT, IMV, IMR, IMP, and IMY, respectively.

Then, I modified the gain and cut-off frequencies of 4th-order eliptic low-pass filters.
Since IMV and IMR has a strong resonance at the observation band, I set notch filters for these resonance at FM8 (notch) of IMV and IMR filter banks.

Theoretically new filters should work but I haven't checked yet because today is a maintainance day for DGS.
So, I wil check them after finishing today's maintanance.

Images attached to this comment
Search Help
×

Warning

×